# Topology-Enhanced Superconducting Qubit Networks for In-Sensor Quantum Information Processing

### v1

@@ 7 From Detection to Information Processing: Flux-Qubit Arrays as Quantum Reservoirs
$f_{k}(t)=\mathcal{A}\,\sin\bigl(\omega_{k}\,t\bigr{missing}),\qquad\omega_{k}=\omega_{\min}+s_{k}\,(\omega_{\max}-\omega_{\min}),$
where $\mathcal{A}$ is a fixed amplitude and $\omega_{k}$ varies linearly between $\omega_{\min}$ and $\omega_{\max}$ as $s_{k}$ runs from 0 to 1. For each input $s_{k}$, the flux‐qubit network is initialized in the ground state of the unperturbed Hamiltonian $H_{0}$, denoted by |$\ket{0}$. The chosen set of observables $\{\hat{O}_{\alpha}\}$ (as in Eq. (23) ) are accumulated in a time‐multiplexed scheme at $n_{t}$ equally spaced instants
@@
where $\gamma\in(0,1)$ controls the memory decay, $\hat{S}^{\,n_{S}}$ is the cyclic shift operator of dimension $l_{r}$ that permutes each component forward by $n_{S}$ positions, $[\hat{S}^{\,n_{S}}]_{ij}=\delta_{(i+n_{S})\bmod l_{r},\,j},$ and $\hat{B}^{\,l_{r}}$ is a lengthening operator that embeds $\vec{m}_{k}$ into an $l_{r}$-dimensional vector by interleaving zeros, ensuring that $\vec{r}_{k}$ has fixed dimension $l_{r}$. The reservoir state is initialized as $\vec{r}_{0}=\mathbf{0}$.
@@
Figure 10: Expectation values $\langle\sigma_{z}^{(i)}\rangle$ for each of the five qubits as a function of the drive frequency $\omega/\omega_{0}$, evaluated at a fixed time $t=2\pi/\omega_{\min}$ for LA configuration (left panel) and CA configuration (right panel). Parameters: normalized flux $f=\Phi_{\mathrm{ex}}/\Phi_{0}=0.45$; inhomogeneous tunneling energies $\Delta_{i}$ linearly distributed in $\Delta[1-\delta,1+\delta]$, with $\delta=0.1$; inductive coupling energy $-MI_{S}^{2}=0.2\,\hbar\omega_{0}$, with $\hbar\omega_{0}=I_{S}\Phi_{0}$.

### v2

@@ 8 From Detection to Information Processing: Flux-Qubit Arrays as Quantum Reservoirs
$f_{k}(t)=\mathcal{A}\,\sin\bigl(\omega_{k}\,t\bigr.),\qquad\omega_{k}=\omega_{\min}+s_{k}\,(\omega_{\max}-\omega_{\min}),$
[…]
@@
where $\gamma\in(0,1)$ controls the memory decay, $\hat{S}^{\,n_{S}}$ is the cyclic shift operator of dimension $l_{r}$ that permutes each component forward by $n_{S}$ positions, $\hat{S}^{\,n_{S}}]_{ij}=\delta_{(i+n_{S})\bmod l_{r},\,j},$ and $\hat{B}^{\,l_{r}}$ is a lengthening operator that embeds $\vec{m}_{k}$ into an $l_{r}$-dimensional vector by interleaving zeros, ensuring that $\vec{r}_{k}$ has fixed dimension $l_{r}$. The reservoir state is initialized as $\vec{r}_{0}=\mathbf{0}$.
@@
Figure 10: Expectation values $\langle\sigma_{z}^{(i)}\rangle$ for each of the five qubits as a function of the drive frequency $\omega/\omega_{0}$, evaluated at a fixed time $t=2\pi/\omega_{\min}$ for LA configuration (left panel) and CA configuration (right panel). Parameters: normalized flux $f=\Phi_{\mathrm{ex}}/\Phi_{0}=0.45$; inhomogeneous tunneling energies $\Delta_{i}$ linearly distributed in $\Delta[1-\delta,1+\delta]$, with $\delta=0.1$; inductive coupling energy $MI_{S}^{2}=0.2\,\hbar\omega_{0}$, with $\hbar\omega_{0}=I_{S}\Phi_{0}$.